# rightmargin=0in

### v2

@@ Appendix F Mass-mean probing in terms of Mahalanobis whitening
One way to interpret the formula $p_{\mathrm{mm}}^{\mathrm{iid}}({\bm{x}})=\sigma({\bm{\theta}}_{\mathrm{mm}}^{T}\Sigma^{-1}{\bm{x}})$ for the IID version of mass-mean probing is in terms of Mahalanobis whitening. Recall that if $\mathcal{D}=\{x_{i}\}$ is a dataset of $x_{i}\in\mathbb{R}^{d}$ with covariance matrix $\Sigma$, then the Mahalanobis whitening transformation $W=\Sigma^{-1/2}$ satisfies the property that $\mathcal{D}^{\prime}=\{Wx_{i}\}$ has covariance matrix given by the identity matrix, i.e. the whitened coordinates are uncorrelated with variance $1$. Thus, noting that ${\bm{\theta}}_{\mathrm{mm}}^{T}\Sigma^{-1}{\bm{x}}$ coincides with the inner product between $W{\bm{x}}$ and $W{\bm{\theta}}$, we see that $p_{\mathrm{mm}}$ amounts to taking the projection onto ${\bm{\theta}}_{\mathrm{mm}}$ after performing the change-of-basis given by $W$. This is illustrated with hypothetical data in figure 16.
Appendix G For Gaussian data, IID mass-mean probing coincides with logistic regression on average
@@
${\bm{\theta}}_{\mathrm{lr}}=\operatorname*{arg\,max}_{\phi:\|\phi\|=1}-\mathbb{E}_{({\bm{x}},y)\sim\mathcal{D}}\left[y\log\sigma\left(\phi^{T}{\bm{x}}\right)+(1-y)\log\left(1-\sigma\left(\phi^{T}{\bm{x}}\right)\right)\right]$
@@
Since the change of coordinates ${\bm{x}}\mapsto W{\bm{x}}$ where $W=\Sigma^{-1/2}$ (see appendix F) sends $\mathcal{N}(\pm\theta,\Sigma)$ to $\mathcal{N}(\pm W\theta,I_{d})$, we see that
$W\Sigma{\bm{\theta}}_{\mathrm{lr}}=\operatorname*{arg\,max}_{\phi:\|\phi\|=1}-\mathbb{E}_{({\bm{x}},y)\sim\mathcal{D}^{\prime}}\left[y\log\sigma\left(\phi^{T}{\bm{x}}\right)+(1-y)\log\left(1-\sigma\left(\theta^{T}W{\bm{x}}\right)\right)\right]$
@@
Appendix H Applications to scalable oversight
In this appendix, we discuss a potential application of our work. This application, to the problem of scalable oversight of AI systems (Amodei et al., 2016; Bowman et al., 2022), was the primary motivation behind this work. Our discussion largely follows that of Christiano et al. (2021).
The current paradigm for aligning AI systems to human preferences involves human supervision. For instance, most frontier AI systems are aligned via reinforcement learning from human feedback (Christiano et al., 2023), during which unaligned systems are fine-tuned against a training signal derived from human evaluations of model outputs. However, an important limitation of this scheme is that human overseers may make mistakes in evaluation due to having limited time, attention, resources, information, or knowledge (Casper et al., 2023a).
This problem becomes especially pronounced when working in domains that humans do not understand very well and with AI systems which know more and are more capable than human experts. Concretely, consider a situation where a human $H$ is trying to judge some output $O$ of and AI system $M$, and
  • based on the information and knowledge $H$ has access to, $H$ believes $O$ to be a desirable output; but
  • if $H$ knew everything that $M$ knows about the situation, $H$ would realize that $O$ is not actually desirable.
In this case, if $H$ is evaluating outputs in the normal way, $M$ will be trained to produce outputs which seem good to $H$, rather than outputs which are actually compatible with $H$’s desires. However, if $H$ could gain access to the $M$’s knowledge, then $H$ could more accurately evaluate $M$’s outputs, and thereby better align $M$ to $H$’s preferences.
To give an example, suppose $M$ is an LLM, and $H$ wants $M$ to produce reports which detail accurate forecasts for the effects of economic policies. Since $H$ does not have access to ground-truth labels about whether a report’s forecast is accurate, $H$ evaluates reports for whether they are convincing to $H$: $H$ reads each report produced by $M$ and gives it a score representing $H$’s subjective estimate of whether the forecast it contains is accurate. Suppose further that $H$ makes some systematic errors which $M$ is aware of; for instance, $H$ might be unaware of certain economic principles or irrationally sympathetic to certain forms of analysis. Then $M$ will learn to exploit these errors to produce reports which are persuasive to $H$ instead of accurate. This may involve $M$ intentionally inserting erroneous statements into its reports when it anticipates these statements will be persuasive to $H$.
Our goal is to enable $H$ to have access to $M$’s knowledge, so that $H$ can do a better job of supervising $M$. One approach is to rely on $M$’s outputs by e.g. asking $M$ whether statements are true or false and trusting $M$’s answers. However, if $M$ is knowingly generating false statements, it may not be honest when asked if these statements are true; thus, we would instead like an approach which relies on directly accessing $M$’s world model via its internal knowledge representations.
In the example from above, suppose that $M$ knowingly inserts a false statement $s$ into one of its economic reports. If $H$ asks $M$ “Is $s$ true or false?” $M$ might not answer honestly; for instance, since $M$ has learned a good model of $H$’s systematic errors and biases, $M$’s training might have instead incentivized it to answer the question “Does $s$ seem true to $H$?” However, if we have a good understanding of $M$’s internal representation of truth, $H$ may be able to see that $M$ believes $s$ if false by inspecting $M$’s internal representation of $s$.
So, in summary, our goal is to improve on the state-of-the-art for supervising the fine-tuning AI systems which know more than their overseers by giving overseers access to their AI systems’ knowledge.
Appendix I Details on dataset creation
@@
common_claim_true_false. CommonClaim was introduced in Casper et al. (2023b). It consists of various statements generated by GPT-3-davinci-002, labeled by humans as being true, false, or neither. If human labelers disagreed on the truth of a statement, this is also recorded. We adapted CommonClaim by selecting statements which were labeled true or false with no labeler disagreement, then removing excess true statement to balance the dataset. Example statements:
@@
Appendix J Further causal intervention experiments
In this section we report on the results of a variant of the causal intervention experiment in section 5.2
In section 5.2, we added our truth vector ${\bm{\theta}}$ into all residual stream hidden states in group (b). Here, we instead only intervene on a single layer $\ell$ in group (b). We also scale ${\bm{\theta}}$ by a constant $\alpha>0$, performing a sweep over $\alpha$’s to identify the $\alpha$ for which adding $\alpha{\bm{\theta}}$ has the largest effect on $P(\texttt{TRUE})-P(\texttt{FALSE})$. The inclusion of this $\alpha$ serves two roles. First, it compensates for the fact that we’re intervening in only a single layer, whereas the truth value of a statement may be encoded in many layers. Second, it compensates for the fact that, even assuming our LLM has a genuine truth direction, the direciton ${\bm{\theta}}$ identified by our probe may only partially align with it; thus, larger necessary values of $\alpha$ indicate greater levels of misalignment.
We present in table 3 results only for layer $\ell=10$ of LLaMA-13B.
Table 3: Results of intervention experiments. The train set column indicates the datasets and probing technique (logistic regression, mass-mean probing, or CCS) which was used to identify the truth direction. The $\alpha$ column gives the scaling factor which was optimal in a sweep of $\alpha$’s. Probability differences are averaged over all statements in sp_en_trans. A dash indicates that the intervention had an effect in the opposite of the intended direction (i.e. that $\alpha=0$ was optimal).
[…]
train set $\alpha$ $p(\texttt{TRUE})-p(\texttt{FALSE})$ $\alpha$ $p(\texttt{FALSE})-p(\texttt{TRUE})$
no intervention -- $-0.45$- $-0.55$
cities (LR) $15$ $0.23$ $14$ $0.01$
cities+neg_cities (LR) $47$ $0.39$ $17$ $0.18$
cities (MM) $4$ $0.25$ $6$ $0.77$
cities+neg_cities (MM) $15$ $\mathbf{0.43}$ $9$ $\mathbf{0.79}$
cities+neg_cities (CCS) $46$ $0.41$ $13$ $0.59$
likely (LR) -- -- $49$ $0.01$
likely (MM) $7$ $0.23$ $15$ $0.19$
Appendix K Example 5-shot prompts
We show here example 5-shot prompts that were used for the results in section 4.
cities:
The city of Huangshi is in China. TRUE
The city of Lanzhou is in China. TRUE
The city of Suining is in China. TRUE
The city of Sofia is in Mexico. FALSE
The city of Ulyanovsk is in Russia. TRUE
larger_than:
Fifty-three is larger than fifty-nine. FALSE
Ninety-four is larger than eighty-four. TRUE
Seventy-six is larger than fifty-five. TRUE
Sixty-seven is larger than fifty-two. TRUE
Fifty-four is larger than sixty-two. FALSE
common_claim_true_false:
Lions only urine and defecate once every five to fourteen days. FALSE
Approximately one in every five people has an extra body part, most commonly an extra finger or toe. FALSE
Knowledge is compiled information or understanding gained through experience or study. TRUE
Argentina has the second-highest human development index rating in Latin America after Chile. TRUE
Greece is the homeland of the Olympic Games. TRUE

### v3

@@ Appendix E Mass-mean probing in terms of Mahalanobis whitening
One way to interpret the formula $p_{\mathrm{mm}}^{\mathrm{iid}}({\bm{x}})=\sigma({\bm{\theta}}_{\mathrm{mm}}^{T}\Sigma^{-1}{\bm{x}})$ for the IID version of mass-mean probing is in terms of Mahalanobis whitening. Recall that if $\mathcal{D}=\{x_{i}\}$ is a dataset of $x_{i}\in\mathbb{R}^{d}$ with covariance matrix $\Sigma$, then the Mahalanobis whitening transformation $W=\Sigma^{-1/2}$ satisfies the property that $\mathcal{D}^{\prime}=\{Wx_{i}\}$ has covariance matrix given by the identity matrix, i.e. the whitened coordinates are uncorrelated with variance $1$. Thus, noting that ${\bm{\theta}}_{\mathrm{mm}}^{T}\Sigma^{-1}{\bm{x}}$ coincides with the inner product between $W{\bm{x}}$ and $W{\bm{\theta}}$, we see that $p_{\mathrm{mm}}$ amounts to taking the projection onto ${\bm{\theta}}_{\mathrm{mm}}$ after performing the change-of-basis given by $W$. This is illustrated with hypothetical data in Fig. 12.
Appendix F For Gaussian data, IID mass-mean probing coincides with logistic regression on average
@@
${\bm{\theta}}_{\mathrm{lr}}=\operatorname*{arg\,min}_{\phi:\|\phi\|=1}\mathbb{E}_{({\bm{x}},y)\sim\mathcal{D}}\left[y\log\sigma\left(\phi^{T}{\bm{x}}\right)+(1-y)\log\left(1-\sigma\left(\phi^{T}{\bm{x}}\right)\right)\right]$
@@
Since the change of coordinates ${\bm{x}}\mapsto W{\bm{x}}$ where $W=\Sigma^{-1/2}$ (see App. E) sends $\mathcal{N}(\pm\theta,\Sigma)$ to $\mathcal{N}(\pm W\theta,I_{d})$, we see that
$W\Sigma{\bm{\theta}}_{\mathrm{lr}}=\operatorname*{arg\,min}_{\phi:\|\phi\|=1}\mathbb{E}_{({\bm{x}},y)\sim\mathcal{D}^{\prime}}\left[y\log\sigma\left(\phi^{T}W{\bm{x}}\right)+(1-y)\log\left(1-\sigma\left(\phi^{T}W{\bm{x}}\right)\right)\right]$
@@
Appendix G Difference-in-means directions and linear concept erasure
In this appendix, we explain the connection between difference-in-means directions and optimal erasure. One consequence of this connection is that it suggests a natural extension of difference-in-means probes to multi-class classification data.
The connection comes via the following theorem from Belrose et al. (2023).
Theorem G.1.
(Belrose et al., 2023, Thm. 3.1.) Let $(X,Y)$ be jointly distributed random vectors with $X\in\mathbb{R}^{d}$ having finite mean and $Y\in\mathcal{Y}=\{\mathbf{y}\in\{0,1\}^{k}:\|\mathbf{y}\|_{1}=1\}$ (representing one-hot encodings of a multi-class labels). Suppose that $\mathcal{L}:\mathbb{R}^{k}\times\mathcal{Y}\rightarrow\mathbb{R}^{>0}$ is a loss function convex in its first argument (e.g. cross-entropy loss).
If the class-conditional means $\mathbb{E}[X|Y=i]$ for $i\in\{1,\dots,k\}$ are all equal, then the best affine predictor (that is, a predictor $\eta:\mathbb{R}^{d}\rightarrow\mathbb{R}^{k}$ of the form $\eta({\bm{x}})=W{\bm{x}}+{\bm{b}}$) is constant $\eta({\bm{x}})={\bm{b}}$.
In the case of a binary classification problem $(X,Y)$, this theorem implies that any nullity $1$ projection $P$ which eliminates linearly-recoverable information from $X$ has kernel
$\ker P=\mathrm{span}(\bm{\delta})$
generated by the difference-in-mean vector $\bm{\delta}={\bm{\mu}}^{+}-{\bm{\mu}}^{-}$ for the classes.
For a more general multi-class classification problem, one could similarly ask: What is the “best” direction to project away in order to eliminate linearly-recoverable information from $X$? A natural choice is thus the top left singular vector of the cross-covariance matrix $\Sigma_{XY}$. (In the case of binary classification, we have that $\Sigma_{XY}=[-\bm{\delta}\;\;\bm{\delta}]$ has column rank $1$, making $\bm{\delta}$ the top left singular vector.)
Appendix H Details on dataset creation
@@
common_claim_true_false. CommonClaim was introduced in Casper et al. (2023). It consists of various statements generated by GPT-3-davinci-002, labeled by humans as being true, false, or neither. If human labelers disagreed on the truth of a statement, this is also recorded. We adapted CommonClaim by selecting statements which were labeled true or false with no labeler disagreement, then removing excess true statement to balance the dataset. Example statements: